DX LUCI2 LBTB AO Science

01:10 Closed loop
  • 10.2 mag catalog
  • 12.4 measured
  • Bin 2, 400 Hz
  • Gopt = 0.23
  • NCPA On
  • Seeing 0.97->1.2

01:25 Skip Frame, loop open
  • Paused LUCI2
  • Re-closed loop and set gains
  • Not clear what the cause was

01:45 Skip Frame, loop open
  • Saw a rolled edge in Command and Current at 11:30->12:00 when offset was requested
    • Not clear if offset caused it or rolled edge happened just before offset and that shape was then frozen.

03:21 Done with AO for a while. Will restart AO at 23:30 MST or so

07:23 Start AO again
  • 15.4 mag catalog
  • 15.8 mag measured
  • bin 4, 100 Hz
  • 3.4, 3.0, 3.3 from Optimize Gain
  • 1.5, 1.2, 1.2 Set manually with same RMS
  • Gopt would be set to 0.8, or get large number with KO. Gopt Not used
  • NCPA Off

10:31 ESM Preset to new object
  • PresetAO fail at 10:31:54.8 right after Setting tip tilt amp - 0.0
  • [Errno 4] Interrupted system call

10:32 Resent ESM preset. Succeeded
  • 2.0, 1.5, 1.5 Set manually
  • Bin 4, 200 Hz, 12 modes
  • 15 mag measured

12:17 ESM Preset
  • 2.0, 1.5 1.5
  • Bin 4, 400 Hz, 12 modes
  • 12.4 mag measured

12:57 ResumeAO failed. Not sure why
  • AdSec left in Failed state
  • Recover failure would not work after adsecarb stop/start
  • Did an IDL and Fastdiagn process stop/start but still get error trying to SetFlatAO
  • Did a full adsc_stop/start. Jochen send new preset. All ok
  • New preset at 13:04
  • 13:23 Loop closed. Had trouble with pupil centering. Had to do it manually. Not sure why

13:45 Background light getting high. Stop observing

-- %USERSIG{DougMiller - 2017-12-31}%
Topic revision: r5 - 31 Dec 2017, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback